# iKap: Kinematics-aware Planning with Imperative Learning

### v1

@@ IV-B4 Real-time Demo on Real Robot
To evaluate the real-time online planning capabilities of iKap in real-world environments, we conduct tests using the quadruped robot Unitree GO2 equipped with an Intel RealSense D435i camera for depth perception. The robot operates with bicycle kinematics model, featuring a minimum turning radius of 0.73 meter. The depth measurements are inherently noisy and can suffer from distribution shifts compared to simulation data. Despite these challenges, our planning network, although trained exclusively in simulation without fine-tuning on real-world data, demonstrates strong generalization to real robot deployment. As shown in Fig. 6, the robot can conduct challenging navigation tasks where the robot starts indoors, passes through a narrow corridor, circles around a public area, and returns to the room along the same path. Additionally, we command the robot to go from off-road to on-road with a large stone blocking it. During this process, the human operator is responsible for sending goal points, while local navigation is autonomously managed by the planner.
@@ V Conclusion and Limitation
We present iKap, a novel planning system that integrates kinematics into an end-to-end learning pipeline for path planning and control. By combining depth perception with a differentiable MPC module, iKap predicts dynamically feasible waypoints while ensuring that the planned trajectories comply with the robot’s kinematic constraints. Experimental results demonstrate that iKap enhances success rates in complex simulated environments and operates effectively under real-world conditions even with noisy depth perception, making it a promising solution for practical robot navigation. However, a limitation of this work is that we have only tested our model in human-made structured environments. In the future, we plan to extend iKap by developing a time-varying cost map for dynamic obstacles and testing the real-world performance under various environments.

### v3

@@ IV-B4 Real-time Demo on Real Robot
To evaluate the real-time online planning capabilities of iKap in real-world environments, we conduct tests using the quadruped robot Unitree GO2 equipped with an Intel RealSense D435i camera for depth perception. The robot operates with bicycle kinematics model, featuring a minimum turning radius of 0.73 meters. The depth measurements are inherently noisy and can suffer from distribution shifts compared to simulation data. Despite these challenges, our planning network, although trained exclusively in simulation without fine-tuning on real-world data, demonstrates strong generalization to real robot deployment. As shown in Fig. 6, the robot can conduct challenging navigation tasks where the robot starts indoors, passes through a narrow corridor, circles around a public area, and returns to the room along the same path. Additionally, we command the robot to go from off-road to on-road with a large stone blocking it. During this process, the human operator is responsible for sending goal points, while local navigation is autonomously managed by the planner.
@@ V Conclusion and Limitation
We present iKap, a novel planning system that integrates kinematics into an end-to-end learning pipeline for path planning and control. By combining depth perception with a differentiable MPC module, iKap predicts waypoints while ensuring that the planned trajectories comply with the robot’s kinematic constraints. Experimental results demonstrate that iKap enhances success rates in complex simulated environments and operates effectively under real-world conditions even with noisy depth perception, making it a promising solution for practical robot navigation. Future work includes evaluating this work across diverse environments beyond human-made structured environments and diverse robot platforms such as Clearpath Husky Observer, as well as extending iKap into three-dimensional space.
Acknowledgements
This work was partially supported by the DARPA award HR00112490426 and ONR award N00014-24-1-2003. Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the authors and do not necessarily reflect the views of DARPA or ONR.